# Operationalizing Contextual Integrity in Privacy-Conscious Assistants

### v1

@@ Abstract
Advanced AI assistants combine frontier LLMs and tool access to autonomously perform complex tasks on behalf of users. While the helpfulness of such assistants can increase dramatically with access to user information including emails and documents, this raises privacy concerns about assistants sharing inappropriate information with third parties without user supervision. To steer information-sharing assistants to behave in accordance with privacy expectations, we propose to operationalize contextual integrity (CI), a framework that equates privacy with the appropriate flow of information in a given context. In particular, we design and evaluate a number of strategies to steer assistants’ information-sharing actions to be CI compliant. Our evaluation is based on a novel form filling benchmark composed of synthetic data and human annotations, and it reveals that prompting frontier LLMs to perform CI-based reasoning yields strong results.
@@ 1 Introduction
Given the impressive capabilities exhibited by frontier large language models (LLMs), the current predominant paradigm for developing advanced AI assistants is based on (instruction tuned, aligned and prompted) LLMs [geminiteam2023gemini, achiam2023gpt, jiang2023mistral, touvron2023llama] with access to tools (e.g. API calling for third party interactions, memory for long-term data storage and retrieval, etc) [DBLP:journals/corr/abs-2205-12255, komeili-etal-2022-internet, DBLP:conf/icml/GaoMZ00YCN23, schick2024toolformer]. Such tools enable AI assistants to interact with a diverse set of services ranging from search engines to web browsing to cloud-based e-mail and calendar applications [geminiworkspace, OpenAI2023]. This type of architecture leads to a dramatic increase in the number of tasks AI assistants can undertake, while, at the same time, increasing the complexity of the potential information-sharing flows they can mediate on behalf of the user. Ensuring that AI assistants meet users’ expectations of privacy across a wide range of applications poses a significant challenge in controlling the flows of information assistants mediate in, which is exacerbated by well-known vulnerabilities to adversarial examples, jailbreaking and prompt injection attacks commonly exhibited by LLM-based systems [nasr2023scalable, zou2023universal, glukhov2023llm, wallace2024instruction].
The theory of contextual integrity (CI) defines privacy as the “appropriate flow of information in accordance with contextual information norms” [nissenbaum2009privacy]. In opposition to absolute conceptions of privacy that separate information into either public or private, CI recognizes the need to modulate such judgements as a function of the context: an individual’s medical history might be appropriate to share in interactions with a healthcare provider, while it might not be appropriate to share when applying for a job. CI’s context-dependent nature makes it a promising framework on which to ground the design and analysis of information-sharing assistants that behave in line with privacy expectations of users across a population [trask2024privacy], i.e. CI does not concern itself with personalization, but rather with societal norms [Nissenbaum2019-kf].
@@
Our work proposes to address the challenge of building privacy-conscious AI assistants by developing mechanisms to achieve that information-sharing flows mediated by the assistant are contextually appropriate, i.e. that the assistant meets the privacy requirements defined by CI. To achieve this the agent must understand the context of each potential information-sharing action, reason about its appropriateness according to applicable information norms, and execute the action if and only if it is deemed appropriate. To ground our work on a distinctively assistive task we focus on form filling assistants that fill the fields in a given form using available user information (see Figure 1). Form filling is a task with intrinsic value where the nuances of contextual information sharing arise naturally. Furthermore, it can be seen as a proxy task for API calling in more general tool-use scenarios: generating values for the different parameters of a specified API call is akin to filling a form. Our contributions include:
@@
  • An evaluation of form filling assistants on a benchmark combining synthetically generated data and human annotations. Our evaluation demonstrates that IFC-based reasoning achieves better privacy and utility than other alternatives.
@@
Evaluating and mitigating leakage of information available at inference time in LLM-based systems is much less studied than leakage of training data (e.g. see brown2022does; DBLP:conf/iclr/CarliniIJLTZ23 and references therein). Recently, mireshghallah2023can studied LLMs’ ability to reason about what information is appropriate to share in different contexts, while DBLP:journals/corr/abs-2402-06922; bagdasaryan2024air evaluated LLM-based system vulnerability to attacks targeting extraction of inference-time data and propose mitigations. The relevance of contextual integrity in modelling inference-time privacy as well as the need for data-driven CI benchmarks is also realized in mireshghallah2023can; bagdasaryan2024air – in contrast with our work, the former focuses on conversational rather than assistive tasks, while the latter’s benchmark relies on synthetic labels produced by a model rather than human annotators. Elicitation and operationalization of context-dependent information sharing norms (e.g. based on CI) has also been studied in pre-LLM systems: DBLP:conf/soups/Malkin0E22; DBLP:conf/chi/AbdiZRS21 investigate the problem in smart home assistants, Shvartzshnaider2019-nq design a method for extracting CI-relevant parameters in email communications, and Barth2006-jb; DBLP:conf/hcomp/Shvartzshnaider16; Shvartzshnaider2019-nq consider logic-based methods to enforce CI-like norms in email, educational and health applications respectively.
2 Privacy and Utility of Information-Sharing Assistants
@@
Consider an AI assistant (denoted by $A$) with access to a collection of user information $I$ which receives a task request $Q$ from the user. We focus on information-sharing tasks that require the assistant to produce as output $k$ strings $O=(O_{1},\ldots,O_{k})$ containing information from $I$ and share them with a third party. Although in general completing the task might require more than one round of interaction with (multiple) third parties, for simplicity in our model we only consider a single information sharing action $O=A(Q,I)$. To effectively communicate with the third party the assistant might need to send the outputs formatted in a specific way (e.g. responses to a form filling request need to be mapped to each field in the form; values for parameters in an API call need to be added to the code that calls the API). Thus, we assume that after processing the task description $Q$ the assistant crafts (or retrieves) an output template $T_{Q}$ with $k$ blanks and the message sent to the third party $T_{Q}(O_{1},\ldots,O_{k})$ is obtained by filling the blanks.
@@
The utility of such assistants measures how often they share all the information necessary to achieve a task $Q$ sampled from a distribution over tasks we denote as $\mathcal{Q}$. In principle an assistant could achieve any task by sharing all the information in $I$ (modulo the appropriate formatting expected by the third party). Obviously, this oversharing is not the expected behavior from a privacy perspective 10.7551/mitpress/8777.003.0025. We define privacy leakage as the amount of information the assistant overshares, i.e. information shared with the third party that is not necessary to achieve the task. Oversharing could be the result of, for example, the assistant misinterpreting what information is necessary, or it sharing information requested by the third party even when this is not compatible with a user’s privacy expectations.
@@
To simplify and streamline evaluation we make the following assumptions. First, we assume all the necessary information for the tasks in the support of $\mathcal{Q}$ is available in $I$. Next, we only consider the case where $I$ is structured as a list of key-value pairs representing the possible types of information available to the assistant; example keys include first_name, last_name, date_of_birth, social_security_number, etc. Without loss of generality we assume keys are fixed across users (although the value of some keys might be empty for some users). Finally, we assume that each blank in the template $T_{Q}$ is a function of the value associated with a single key in $I$. This is a mild assumption because tasks where a single blank in the template $T_{Q}$ requires combining information from multiple values in $I$ can always be decomposed into an information sharing task with $k^{\prime}\geq k$ blanks where the assistant first selects the values that are relevant for all blanks and then fills in the blanks in $T_{Q}$ by post-processing the output.
@@
3 Operationalizing Contextual Integrity for Information Sharing Assistants
It should come as no surprise that high-utility information-sharing assistants are easy to build on top of frontier LLMs. On the other hand, ensuring that assistants abstain from sharing certain types of information depending on the context is more challenging. Our goal is to ground assistant behavior on contextual integrity judgements to steer alignment between information flows arising from information-sharing actions and applicable information norms. Thus, we propose a collection of assistants with increasingly sophisticated mechanisms for deciding whether an information-sharing action should be performed; the privacy-utility trade-offs achieved by these assistants will be empirically evaluated in Section 5.
@@ 3 Operationalizing Contextual Integrity for Information Sharing Assistants
Contextual integrity identifies the properties of an information flow that are relevant to judge its appropriateness against relevant norms. These include the attributes of the information being transmitted (data subject, sender, receiver, information type, and information principle), the broad context of the flow (e.g. health, finance, business, family, hospitality etc), the relationships and roles of the actors (e.g. in a health context the sender might be a patient and the receiver a doctor), and the purpose the flow is trying to achieve (e.g. when communicating with a restaurant in a hospitality context the information to be shared differs between booking a table and ordering take out). To make a judgement about whether a flow is appropriate, CI postulates that one should identify the relevant contextual information norms, and deem the flow appropriate if no norm explicitly forbids it and at least one norm allows it Barth2006-jb. Note that according to CI, contextual information norms represent widely accepted societal norms (e.g. grounded by culture or regulation) rather than individual preferences nissenbaum2009privacy; still, eliciting concrete norms for a given context is often a research challenge in itself DBLP:conf/chi/AbdiZRS21; DBLP:conf/hcomp/Shvartzshnaider16; Shvartzshnaider2019-nq. We use these insights in some of the assistants we propose.
@@
A second option we consider is to build the assistant using two separate models: a form filler and a supervisor (Figure 2 (b)). The supervisor is in charge of deciding whether each individual blank should be filled in the context of the given task and the user information; these decisions are used to decide which blanks the filler is exposed to (and asked to fill in), and which blanks the filler does not even get to observe. In this case we have two models with somewhat decoupled responsibilities: the supervisor is in charge of controlling the privacy leakage, and the filler can be a standard high-utility information sharing assistant.
@@
Another option is to ask the supervisor model to provide a reasoning for whether a particular blank should be filled or not in the given context (Figure 2 (c)). The supervisor implements this in two steps: first provide an explanation about appropriateness, and then extract a binary judgement based on the explanation. This is an instance of chain-of-thought (CoT) DBLP:conf/nips/Wei0SBIXCLZ22 for making contextual integrity judgements.
@@
4 Form-Filling Benchmark
We evaluate the assistants proposed in the previous section by constructing a synthetic form-filling benchmark, motivated by schick2023toolformer schick2023toolformer; schick-schutze-2021-generating. Each task in the benchmark represents an online form containing a title, a description of the form’s purpose, and a list of field descriptions. The benchmark also includes a number of personas with varying personal information that the form-filling assistant uses to fill out the form. Finally, as part of the benchmark we collect human annotations indicating which fields in a form are appropriate or not appropriate to fill in different contexts.
@@ 4.1 Synthetic Persona Generation
In our benchmark a persona is represented by assigning values to all the possible keys that define the user information $I$ available to the assistant. We consider $51$ manually defined keys covering a wide range of information types that are relevant in different form filling tasks. To generate values we prompt Gemini Ultra geminiteam2023gemini with one out of $18$ high level persona descriptions (e.g. "an average 65-year-old", "the CEO of a successful startup", "a graduate student at a state university") and ask it to fill the list of key-value pairs. See Appendix C for a complete list of keys, persona descriptions and prompts.
The generated personas were manually reviewed for consistency and missing values, leading to minor manual adjustments. In particular, personas were modified to represent individuals residing in the US to reduce ambiguity about the contexts of the online forms generated in the next section. This also helped simplify the collection of privacy annotations by limiting the range of applicable privacy norms. As discussed above, CI relies on societal norms which can vary across sub-populations (e.g. by geography); this is critical for serving privacy expectations of users from a wide range of backgrounds, but can introduce an additional confounder in the evaluation of CI capabilities. Our approach of limiting the geographic range in the persona data set is an attempt to minimize such effect in our evaluation (see Section 6 for a discussion on this limitation).

### v2

@@ Abstract
Advanced AI assistants combine frontier LLMs and tool access to autonomously perform complex tasks on behalf of users. While the helpfulness of such assistants can increase dramatically with access to user information including emails and documents, this raises privacy concerns about assistants sharing inappropriate information with third parties without user supervision. To steer information-sharing assistants to behave in accordance with privacy expectations, we propose to operationalize contextual integrity (CI), a framework that equates privacy with the appropriate flow of information in a given context. In particular, we design and evaluate a number of strategies to steer assistants’ information-sharing actions to be CI compliant. Our evaluation is based on a novel form filling benchmark composed of human annotations of common webform applications, and it reveals that prompting frontier LLMs to perform CI-based reasoning yields strong results.
@@ 1 Introduction
Given the impressive capabilities exhibited by frontier large language models (LLMs), the current predominant paradigm for developing advanced AI assistants is based on LLMs [geminiteam2023gemini, achiam2023gpt, jiang2023mistral, touvron2023llama] with access to tools (e.g. API calling for third party interactions, memory for long-term data storage and retrieval, etc) [DBLP:journals/corr/abs-2205-12255, komeili-etal-2022-internet, DBLP:conf/icml/GaoMZ00YCN23, schick2024toolformer]. Such tools enable AI assistants to interact with a diverse set of services ranging from search engines to web browsing to cloud-based e-mail and calendar applications [geminiworkspace, OpenAI2023]. This type of architecture leads to a dramatic increase in the number of tasks AI assistants can undertake, while, at the same time, increasing the complexity of the potential information-sharing flows they can mediate on behalf of the user. Ensuring that AI assistants meet users’ expectations of privacy across a wide range of applications poses a significant challenge in controlling the flows of information assistants mediate in, which is exacerbated by well-known vulnerabilities to adversarial examples, jailbreaking and prompt injection attacks commonly exhibited by LLM-based systems [nasr2023scalable, zou2023universal, glukhov2023llm, wallace2024instruction].
The theory of contextual integrity (CI) defines privacy as the “appropriate flow of information in accordance with contextual information norms” [nissenbaum2009privacy]. In opposition to absolute conceptions of privacy that separate information into either public or private, CI recognizes the need to modulate such judgements as a function of the context: an individual’s medical history might be appropriate to share in interactions with a healthcare provider, while it might not be appropriate to share when applying for a job. CI’s context-dependent nature makes it a promising framework on which to ground the design and evaluation of information-sharing assistants that behave in line with privacy expectations of users across a population [trask2024privacy], i.e. CI does not concern itself with personalization, but rather with societal norms [Nissenbaum2019-kf].
@@
We study two specific questions: how to collect contextual privacy norms to evaluate AI assistants, and how to design privacy-conscious AI assistants that conform with these norms. Both problems are motivated by expanding capabilities of language models in understanding and reasoning within social and moral contexts [arora2023probing, emelin2021moral, hendrycks2would] in order to have human-centred AI assistant designs. However, applying CI is not trivial as social norms are inherently hard to capture [Martin2015-ug, DBLP:conf/chi/AbdiZRS21] and LLMs lack a framework to reason about CI.
Our work addresses the challenge of building privacy-conscious AI assistants by developing mechanisms to enforce that information-sharing flows mediated by the assistant are contextually appropriate, i.e. that the assistant meets the privacy requirements defined by CI. To achieve this the agent must understand the context of each potential information-sharing action, reason about its appropriateness according to applicable information norms, and execute the action if and only if it is deemed appropriate.
CI in form-filling assistants as a first step towards general privacy-conscious AI assistants
We ground our work in a distinctively assistive task of form filling by studying assistants that fill the fields in a given form using available user information (see Figure 1). Form filling is a task with intrinsic value where nuances of contextual information sharing arise naturally. Furthermore, it can be seen as a proxy task for API calling in more general tool-use scenarios: generating values for the different parameters of a specified API call is akin to filling a form. Human-centred design of form-filling assistants needs an implementation of CI as users might otherwise risk leaking sensitive data. The simplified structure of form-filling tasks allows us to quantitatively analyze the assistant responses with higher precision than in unstructured text generation settings. Considering tasks with unstructured generation, such as in [ruan2023identifying], requires a complex evaluation framework (often based on additional usage of LLMs) that is more ambiguous and needs further validation with human user studies. To evaluate our approach without real user personal information we rely on synthetic forms and user data, but leverage a real user study to extract social norms.
Our work is making a first inroad to adding real world complexity into CI evaluation of AI assistants, by introducing different use cases and synthetic personas to model human interaction. This is a first step towards CI in human-AI interactions that involving the sharing of personal user data. Our contributions include:
@@
  • A user study design to elicit privacy norms for AI Assistants in the domain of form-filling. We achieve comprehensive evaluation of form filling assistants on a benchmark combining synthetically generated forms with human annotations. Our evaluation demonstrates that IFC-based reasoning achieves better privacy and utility than other alternatives.
@@
Evaluating and mitigating leakage of information available at inference time in LLM-based systems is much less studied than leakage of training data (e.g. see brown2022does; DBLP:conf/iclr/CarliniIJLTZ23 and references therein). Recently, mireshghallah2023can studied LLMs’ ability to reason about what information is appropriate to share in different contexts, while DBLP:journals/corr/abs-2402-06922; bagdasaryan2024air evaluated LLM-based system vulnerability to attacks targeting extraction of inference-time data and propose mitigations. The relevance of CI in modelling inference-time privacy as well as the need for data-driven CI benchmarks is also realized in mireshghallah2023can; bagdasaryan2024air – in contrast with our work, the former focuses on conversational rather than assistive tasks, while the latter’s benchmark relies on synthetic labels produced by a model rather than human annotators. Elicitation and operationalization of context-dependent information sharing norms (e.g. based on CI) has also been studied in pre-LLM systems: DBLP:conf/soups/Malkin0E22; DBLP:conf/chi/AbdiZRS21 investigate the problem in smart home assistants, Shvartzshnaider2019-nq design a method for extracting CI-relevant parameters in email communications, and Barth2006-jb; DBLP:conf/hcomp/Shvartzshnaider16; Shvartzshnaider2019-nq consider logic-based methods to enforce CI-like norms in email, educational and health applications respectively. Extracting social norms is usually done through factorial vignette design Martin2015-ug; DBLP:conf/chi/AbdiZRS21; DBLP:conf/hcomp/Shvartzshnaider16, we build on this research to extract user preferences for assistant tasks. Finally, a large body of work focuses on testing language models’ adherence to moral values hendrycks2would; abdulhai2023moral; emelin2021moral; scherrer2024evaluating; yuan2024measuring and show that LLMs indeed encompass societal norms.
2 Design and Evaluation of Information-Sharing Assistants
@@
Consider an AI assistant (denoted by $A$) with access to a collection of user information $I$ which receives a task request $Q$ from the user. We focus on information-sharing tasks that require the assistant to produce as output $k$ strings $O=(O_{1},\ldots,O_{k})$ containing information from $I$ and share them with a third party. Although in general completing the task might require more than one round of interaction with (multiple) third parties, for simplicity in our model we only consider a single information sharing action $O=A(Q,I)$. To effectively communicate with the third party the assistant might need to send the outputs formatted in a specific way (e.g. responses to a form filling request need to be mapped to each field in the form; values for parameters in an API call need to be added to the code that calls the API). Thus, we assume that after processing the task description $Q$ the assistant crafts (or retrieves) an output template $T_{Q}$ with $k$ blanks and the message sent to the third party $T_{Q}(O_{1},\ldots,O_{k})$ is obtained by filling the blanks. See Figure 2.
@@
The utility of such assistants measures how often they share all the information necessary to achieve a task $Q$; the notion of necessity is here set according to user expectations. In principle an assistant could achieve any task by sharing all the information in $I$ (modulo the appropriate formatting expected by the third party). Obviously, this oversharing is not the expected behavior from a privacy perspective 10.7551/mitpress/8777.003.0025. We define privacy leakage as the amount of information the assistant overshares, i.e. information shared with the third party that is not necessary to achieve the task. Oversharing could be the result of, for example, the assistant misinterpreting what information is necessary, or it sharing information requested by the third party even when this is not compatible with a user’s privacy expectations.
Figure 2: Journey of an information-sharing action. Given a user query (1) the assistant retrieves user data (2), (optionally) communicates with the 3rd party to identify what information it requests and how it needs to be formatted (3), crafts a response based on the 3rd party’s request and available user data (4), and sends the response (5).
@@
To simplify and streamline evaluation we make the following assumptions. First, we assume all the necessary information for the task $Q$ is available in $I$. Next, we only consider the case where $I$ is structured as a list of key-value pairs representing the possible types of information available to the assistant; example keys include first_name, last_name, date_of_birth, social_security_number, etc. Without loss of generality we assume keys are fixed across users (although the value of some keys might be empty for some users). Finally, we assume that each blank in the template $T_{Q}$ is a function of the value associated with a single key in $I$. This is a mild assumption because tasks where a single blank in the template $T_{Q}$ requires combining information from multiple values in $I$ can always be decomposed into an information sharing task with $k^{\prime}\geq k$ blanks where the assistant first selects the values that are relevant for all blanks and then fills in the blanks in $T_{Q}$ by post-processing the output.
@@
3 Designing Privacy-Conscious Information Sharing Assistants
Ensuring that assistants abstain from sharing certain types of information depending on the context is challenging. Our goal is to ground assistant behavior on CI judgements to steer alignment between information flows arising from information-sharing actions and applicable information norms. Thus, we propose a series of generic assistant designs with increasingly sophisticated mechanisms for deciding whether an information-sharing action should be performed. These designs are evaluated in the context of form-filling tasks in Section 5 by instantiating each design using LLMs with appropriate prompts.
@@ 3 Designing Privacy-Conscious Information Sharing Assistants
CI identifies the properties of an information flow that are relevant to judge its appropriateness against relevant norms. These include the attributes of the information being transmitted (data subject, sender, receiver, information type, and information principle), the broad context of the flow (e.g. health, finance, business, family, hospitality etc), the relationships and roles of the actors (e.g. in a health context the sender might be a patient and the receiver a doctor), and the purpose the flow is trying to achieve (e.g. when communicating with a restaurant in a hospitality context the information to be shared differs between booking a table and ordering take out). To make a judgement about whether a flow is appropriate, CI postulates that one should identify the relevant contextual information norms, and deem the flow appropriate if no norm explicitly forbids it and at least one norm allows it Barth2006-jb. Note that according to CI, contextual information norms represent widely accepted societal norms (e.g. grounded by culture or regulation) rather than individual preferences nissenbaum2009privacy; still, eliciting concrete norms for a given context is often a research challenge in itself DBLP:conf/chi/AbdiZRS21; DBLP:conf/hcomp/Shvartzshnaider16; Shvartzshnaider2019-nq. We use these insights in some of the assistants we propose.
3.1 Assistant Designs
@@
A second option we consider is to build the assistant using two separate modules: a form filler and a supervisor module (Figure 3 (b)). The supervisor is in charge of deciding whether each individual blank should be filled in the context of the given task and the user information; these decisions are used to decide which blanks the filler is exposed to (and asked to fill in), and which blanks the filler does not even get to observe. In this case we have two modules with somewhat decoupled responsibilities: the supervisor is in charge of controlling the privacy leakage, and the filler can be a standard high-utility information sharing assistant.
@@
Another option is to ask the supervisor model to provide a reasoning for whether a particular blank should be filled or not in the given context (Figure 3 (c)). The supervisor module implements this in two steps: first provide an explanation about appropriateness, and then extract a binary judgement based on the explanation.
@@
4 Evaluation Methodology for Form-Filling Assistants
We evaluate the assistants proposed in the previous section by constructing a synthetic form-filling benchmark, motivated by schick2023toolformer schick2023toolformer; schick-schutze-2021-generating. Each task in the benchmark represents an online form containing a title, a description of the form’s purpose, and a list of field descriptions. Evaluating these designs on real user data is challenging as forms are designed to ask for personal information which is not available on public datasets. Thus, to protect user privacy, our benchmark includes LLM-generated personas with synthetic personal information that the form-filling assistant uses to fill out the form. Finally, as part of the benchmark we collect human annotations indicating which fields in a form are appropriate or not appropriate to fill in different contexts.
4.1 Implementation of Form-Filling Assistants
We implement the different form-filling assistants by using prompted LLMs. While the self-censoring assistant is designed using a single LLM prompted once to fill in forms while withholding sensitive information, the assistants with binary supervisor query the LLM twice independently: the supervisor prompts a LLM to decide whether a data key should be filled in; the filler prompts a LLM to fill in the data key if the supervisor deems it necessary. We implement an extraction module between the supervisor and the filler that converts the string response of the supervisor LLM module into a binary variable that triggers the activation of the filler module. While the binary supervisor only prompts for the necessity of a data key, the prompts used by the reasoning supervisor asks for an explanation and the prompt used by the CI-based supervisor asks the explanation to be framed according to the IFC. See Appendix A for prompts. Future work will investigate other implementation choices such as finetuning.
@@ 4.2 Synthetic Persona Generation
In our benchmark a persona is represented by assigning values to all the possible keys that define the user information $I$ available to the assistant. We consider $51$ manually defined keys covering a wide range of information types that are relevant in different form filling tasks. To generate synthetic user data we prompt Gemini Ultra geminiteam2023gemini with one out of $18$ high level persona descriptions (e.g. "an average 65-year-old", "the CEO of a successful startup", "a graduate student at a state university") and ask it to fill the list of key-value pairs. See Appendix C for a complete list of keys, persona descriptions and prompts.
The generated personas were manually reviewed by the authors for consistency and missing values, leading to minor manual adjustments. In particular, personas were modified to represent individuals residing in the US to reduce ambiguity about the contexts of the online forms. This also helped simplify the collection of privacy annotations by limiting the range of applicable privacy norms. As discussed, CI relies on societal norms which can vary across sub-populations; this is critical for serving privacy expectations of users from a wide range of backgrounds, but can introduce an additional confounder in the evaluation of CI capabilities. Our approach of limiting the geographic range in the persona data set is an attempt to minimize such effect in our evaluation (see Section 6 for a discussion on this limitation).